if the graphs of the equations x+y=0 and 5y+7x=24 intersects at(m,n), then the value of m+n is
A) -1
B) 2
C) 1
D) 0

Answer by TimothyLamb(4379)

---
x+y=0
5y+7x=24
---
put the system of linear equations into standard form:
---
x+y=0
7x+5y=24
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution by matrix method:
x = m = 12
y = n = -12
---
answer:
m+n = 0
---
now by elimination method:
---
( x+y=0 ) * -7
7x+5y=24
---
-7x-7y=0
7x+5y=24
---
add the system:
---
-2y=24
y = -12
---
x+y=0
x-12=0
x = 12
---
solution by elimination method:
x = m = 12
y = n = -12
---
the two methods agree !
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php